Prove that the sequence {an /n } is convergent where { an } is a bounded sequence.



1
Expert's answer
2022-02-04T11:56:07-0500

Suppose that the sequence "a_n" is bounded. It means that there is "M>0" such that "|a_n|<M" "\forall n". Therefore, "|\frac{a_n}{n}|<\frac{M}{n}". Since "\frac{M}{n}\rightarrow0", "n\rightarrow\infty", "|\frac{a_n}{n}|\rightarrow0", "n\rightarrow\infty."